OpenStudy (anonymous):

factor x^3+5x^2-49x-245

OpenStudy (anonymous):

f(x) = x^3+5x^2-49x-245 f(-5) = (-5)^3+5(-5)^2-49(-5)-245 = 0 So, x+5 is one of the factors. Use long division/synthetic division to find the other factors.

OpenStudy (anonymous):

Ex: f(x)=x³-5x²+49x-245 Make it equal to zero and Group factor x^3 - 5x^2 + 49x - 245 = 0 (x^3 + 49x) - (5x^2 + 245) = 0 x(x^2 + 49) - 5(x^2 + 49) = 0 Factor out (x^2+49) leaving (x^2 + 49)(x - 5) = 0 x^2 = -49 x = x = +7i x = -7i and x = 5 is the only real root

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

that's an example

OpenStudy (amistre64):

:) nice change up there

OpenStudy (anonymous):

me? thnx